OpenStudy (anonymous):

Which statement best describes the graph of x3 - 3x2 - x + 3? It starts down on the left and goes up on the right and intersects the x-axis at x = -1, 2, and 3. It starts down on the left and goes up on the right and intersects the x-axis at x = -1, 1, and 3. It starts up on the left and goes down on the right and intersects the x-axis at x = -1, 2, and 3. It starts up on the left and goes down on the right and intersects the x-axis at x = -1, 1, and 3.

OpenStudy (anonymous):

@jdoe0001 I know it's either B or D however I dont know which

OpenStudy (jdoe0001):

hmmm are you supposed to use a graphing calculator or not?

OpenStudy (anonymous):

Yea I used desmos

OpenStudy (anonymous):

@Jbo11 what do you think

OpenStudy (jdoe0001):

http://fooplot.com/#W3sidHlwZSI6MCwiZXEiOiJ4XjMtM3heMi14KzMiLCJjb2xvciI6IiNEQjE4MTgifSx7InR5cGUiOjEwMDB9XQ-- It starts Down on the left and goes Up on the right and intersects the x-axis at x = -1, 1, and 3.

OpenStudy (anonymous):

Ok and this one A quadratic equation is shown below: x2 + 18x + 76 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)2 = q, where p and q are integers? Add 9 to both sides of the equation Add 5 to both sides of the equation Subtract 5 from both sides of the equation Subtract 9 from both sides of the equation

OpenStudy (jdoe0001):

well... do you know what a "perfect squatre trinomial" is?

OpenStudy (anonymous):

Yes

OpenStudy (jdoe0001):

ok... so let us start by grouping so \(\bf x^2 + 18x + 76 = 0\implies (x^2 + 18x) + 76 = 0 \\ \quad \\ (x^2 + 18x+{\color{red}{ \square ?}}^2) + 76 = 0\) what number do you think gives us a perfect square trinomial in the parenthesized groiup?

OpenStudy (jdoe0001):

keep in mind that the middle term is 18x and the middle term is also 2(x)(?)

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

\(\large \begin{cases} middle\ term\implies &18x\\ middle\ term\implies &2(x)({\color{red}{ \square ?}}) \end{cases}\)

OpenStudy (anonymous):

Ok so what would it be

OpenStudy (jdoe0001):

so.... one could say that \(\bf 18x=2(x)({\color{red}{ \square ?}})\implies ({\color{red}{ \square ?}})=?\)

OpenStudy (jdoe0001):

what do you think the missing fellow might be?

OpenStudy (anonymous):

Oh I think what you mean is completing the square

OpenStudy (anonymous):

I dont know what it is

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@jhonyy9 @Jonnychewy

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@Emily_dreamsbig

OpenStudy (anonymous):

@Jonnychewy